4) You are asked by the CCU nurse to confirm her calculations of the dopamine drip that she is using on her patient. She states that she is using the standard hospital bag which is 800mg of dopamine in 1000ml and that the patient weighs 183 pounds and the physician ordered the rate to be set at 50mcg/kg/min. She states that she currently has the pump set at 85ml/hour. She want you to check her calculations. You perform the calculations and conclude that: (Round to nearest ml)?
Answer #1

800 mg dopamine in 1000 ml

so 1 ml= 800/1000 =0.8 mg

patient weight = 183 pounds=83,0074 kg

prescribed dose=50 mcg/kg/min 50 mcg=0.05 mg

= 0.05*83.0074*60=249.02= 249 mg/hour

if 1 ml contains 0.8 mg of dopamine,

then 249 mg = ? ml (1:x= 0.8:249

1/x=0.8/249

x=249/0.8=311.25)

249mg* 1 ml/0.8 mg= 311.25 ml

so ,prescribed dose =311.25 ml/hour